# Accelerated Ostwald ripening by chemical activity

### v1

@@ IV.1 Small droplet
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R}\left(\Delta\varphi_{\mathrm{s}}-\frac{\lambda}{R}\right)\left(1+\frac{R}{\ell}\right)-\frac{s_{\mathrm{l}}R}{3}.$ (24)
where we defined the excess saturation $\Delta\varphi_{\mathrm{s}}=\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$ and $\lambda$ is the Laplace pressure correction of the sticky volume fraction in the gas; see Sec. II.1. The growth-enhancing factor $1+R/\ell$ arises from the reactions in the dilute phase, whereas the last term, hindering growth, originates from the reactions in the dense phase. Indeed small droplets, $R\ll\lambda/\Delta\varphi_{\mathrm{s}}$, vanish. In the absence of reactions ($\ell\to\infty$, $s_{\mathrm{l}}=0$), Eq. (24) exactly coincides with the passive Lifshitz-Slyozov growth law [24]. We now elaborate on the results for the active case, which are summarized in Fig. 3.
If $s_{\mathrm{l}}$ is sufficiently large, indefinite growth and ripening are prevented. This follows because sticky proteins are lost at a rate proportional to the volume of the droplet, whereas they join from the dilute phase at a rate proportional to the radius. Thus, even for small $R$, large $s_{\mathrm{l}}$ may be sufficient to counteract the ripening. In this case, the radius $R^{*}=(3D\Delta\varphi_{\mathrm{s}}/s_{\mathrm{l}})^{1/2}$ is a stable point of Eq. (24), where we kept leading order in $\lambda/R^{*}\ll 1$ [3]. This implies that smaller droplets will grow and larger ones will shrink until $R^{*}$ is reached. [^5] Note that as droplets grow, the excess saturation $\varphi_{\mathrm{s}}^{\infty}$ decreases (cf. Eq. (10)), lowering further the terminal radius $R^{*}$. Therefore, $R^{*}=(3D\Delta\varphi_{\mathrm{s}}/s_{\mathrm{l}})^{1/2}$ is in fact a transcendental equation for $R^{*}$. This size control has been reported and discussed in depth, e.g., in Ref. [3] and observed in biological systems [4, 5]. In this case, the large droplet regime is inaccessible.
On the other hand, if the above terminal radius is too large, $R^{*}\gg\ell$, then size control does not occur before the droplets reach the $R\gg\ell$ regime we discuss below. This would be the case if the reactions occur too slowly, $s_{\mathrm{l}}<D_{\mathrm{g}}\Delta\varphi_{\mathrm{s}}/\ell^{2}$ (to leading order in $\lambda/\ell\ll 1$). In that case, the second term in Eq. (24) can be neglected. We comment that the additional factor $1+R/\ell$, originating from the reactions in the dilute phase, accelerates the growth of droplets of size $R\lesssim\ell$, though to a subleading extent. Therefore, in total, we recover standard Ostwald ripening (without meaningful acceleration) which continues until the growth dynamics transitions to the large droplet regime, $R\gg\ell$, discussed next.
IV.2 Large droplet
Figure 3: Phase diagram of droplet dynamics as a function of total dense phase reaction rate and droplet size. Reactions in the dilute phase may accelerate Ostwald ripening (cf. Fig. 2), whereas reactions in the dense phase hinder droplet growth and ripening by converting particles from sticky to nonsticky. Small droplets with $R\ll\lambda/\Delta\varphi_{\mathrm{s}}$ always tend to shrink due to large Laplace pressure. Other regions are approximately bounded by the inter-reaction diffusion length $\ell$. For very fast dense phase reactions, $s_{\mathrm{l}}\gg D_{\mathrm{g}}\Delta\varphi_{\mathrm{s}}/\ell^{2}$, size control is inevitable. Otherwise, droplets can reach the regime $R\gg\ell$, where: For intermediate dense phase reaction rate, $(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell\ll s_{\mathrm{l}}\ll D\Delta\varphi_{\mathrm{s}}/\ell^{2}$, a negative effective surface tension prevents small droplets from shrinking. For slow dense phase reaction rate, $s_{\mathrm{l}}\ll(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell$, Ostwald ripening occurs with an accelerated rate. Parameters: $R$, droplet radius; $s_{\mathrm{l}}$, total dense phase reaction rate; $\Delta\varphi_{\mathrm{s}}$, dilute phase volume fraction relative to the coexistence value; $\lambda$, characteristic width of a droplet interface; $\ell$, characteristic distance between reactions; $k_{i\to j}$, dilute-phase switching rates, $D_{\mathrm{g}}$, dilute-phase diffusion constant.
For larger droplets, solving Eq. (21) for $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ in the limit $R/\ell\gg 1$ yields,
$\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$ $=$ $\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)+\frac{s_{\mathrm{l}}}{k_{\mathrm{ns}\to\mathrm{s}}}$
$+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\left[\left(1+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\right)\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}-\frac{2s_{\mathrm{l}}}{k_{\mathrm{s}\to\mathrm{ns}}}\right]\frac{\ell}{R},$
and inserting the result above in Eq. (22) results in
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D^{\mathrm{eff}}}{R}\left(\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}-\frac{\lambda^{\mathrm{eff}}}{R}\right),$ (26)
where [^6] A reactive lengthscale similar to Eq. (19) emerges for the dense phase. To obtain Eqs. (28) and (29), we assumed that it is the same lengthscale $\ell$ in both phases. If these lengths are different, our results do not change qualitatively; rather, $s_{\mathrm{l}}$ will be multiplied by a finite factor in each expression. We address this detail in SI Appendix, Sec. II.
$D^{\mathrm{eff}}$ $=$ $D_{\mathrm{g}}\left(1+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\right),$ (27)
$\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ $=$ $\Delta\varphi_{\mathrm{s}}-\frac{s_{\mathrm{l}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}},$ (28)
$\lambda^{\mathrm{eff}}$ $=$ $\lambda-\frac{2s_{\mathrm{l}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\ell+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}\ell.$ (29)
Notably, the dilute-phase reactions have successfully increased the effective diffusion coefficient by a factor $1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}$. The effective supersaturation (Eq. (28)), however, is smaller than the actual residual saturation $\Delta\varphi_{\mathrm{s}}$ due to net conversion of sticky material to nonsticky inside the droplet. This implies that, if Ostwald ripening occurs, it will stop earlier than in the passive case due to material exhaustion (Eq. (10)).
We proceed to discuss the possible growth outcomes implied by Eq. (26), summarized in Fig. 3, by inspecting the effective Laplace pressure correction coefficient, $\lambda^{\mathrm{eff}}$. Understanding it is intricate: Growth driven by Laplace pressure, as mentioned above, can only occur with a rate proportional to the radius, while reactions on either side of the interface (consuming and creating sticky protein inside and outside the droplet, respectively) occur with rates proportional to the area. Since now $R\gg\ell$, the growth matching condition, Eq. (21), requires the reactions to counteract one another, thus canceling each other. Hence, the effects described below arise from the next-order corrections; see SI Appendix, Sec. II.
When the reactions inside the droplet are sufficiently fast, $s_{\mathrm{l}}\gg(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell$, we may neglect the contribution from the Laplace pressure, the first term in Eq. (29). Thus, temporarily ignoring the third term in Eq. (29), the reactions cause an effective negative Laplace pressure correction coefficient, $\lambda^{\mathrm{eff}}=-2\ell s_{\mathrm{l}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$. As a result, the droplets do not undergo ripening per se, i.e., smaller droplets do not shrink. Instead, all droplets that pass the $R\gg\ell$ threshold are expected to grow until the extensive, stable radius $R^{\dagger}=\lambda^{\mathrm{eff}}/\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}\sim V^{1/3}$ is reached. For this stable radius, the effective supersaturation becomes negative. Thus, now including the third term of Eq. (29), we see that it will further contribute to the already negative $\lambda^{\mathrm{eff}}=-\ell[s_{\mathrm{l}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})+(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})|\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}|]$. As growth continues, material is being taken up from the dilute phase, so $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ becomes even more negative so $|\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}|$ increases. This ripening arrest has been recently reported in Ref. [6]; we will not investigate it here further.
We now discuss the sought-after regime where Ostwald ripening is accelerated. For slow reactions inside the droplet, $s_{\mathrm{l}}\ll(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell$, the Laplace pressure dominates as the second term of Eq. (29) is negligible. Thus, the Laplace pressure correction coefficient is positive, and ripening occurs as expected. Here, $\lambda^{\mathrm{eff}}=\lambda+\ell(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$, where the effective supersaturation, $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$, as in standard Ostwald ripening, is positive. Since $\ell/\lambda$ is a large number, the contribution from the small $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ may be transiently important. During a transient where $\lambda\ll\ell(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$, we find $\mathrm{d}R/\mathrm{d}t=(D^{\mathrm{eff}}\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}/R)[1+(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})(\ell/R)]\simeq D^{\mathrm{eff}}\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}/R$, meaning that $R^{2}\sim t$ until the effective supersaturation lowers sufficiently due to ripening.
With the progression of ripening, once $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ decreases sufficiently, we reproduce the standard $\lambda^{\mathrm{eff}}=\lambda$, and hence also the scaling of Lishitz-Slyozov [24], but with a larger diffusion constant compared to the passive Eq. (11),
$\langle R(t)\rangle^{3}-\langle R(0)\rangle^{3}_{\mathrm{a}}=\frac{4}{9}D^{\mathrm{eff}}\lambda t.$ (30)
Note that due to the transient for $R\ll\ell$ and possibly altered nucleation rates [42, 1], the intercept $\langle R(0)\rangle^{3}_{\mathrm{a}}$ might be different from the one in the passive case. Thus, the average droplet volume grows by a factor $D^{\mathrm{eff}}/D_{\mathrm{g}}=1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}$ faster compared to the passive case. Interestingly, the faster the switching from sticky to nonsticky compared to the reverse process, the faster is the ripening rate. This increase of $D^{\mathrm{eff}}$ occurs at the expense of having a lower sticky supersaturation, Eq. (16). The fact that this is the only regime and mechanism by which Ostwald ripening can be accelerated is the main result of the paper.
IV.3 Realizability of accelerated ripening
With our theoretical results for accelerated ripening laid out, we assess their experimental relevance. Pragmatically, it is not trivial to realize the above constellation of assumptions. For clarity, we recap our approximations and explain the sacrifices made so as to achieve acceleration. Despite these shortcomings, we will argue for the feasibility to accelerate ripening and present the experimental results of Ref. [15] as evidence for a workable protocol.
The first pitfall is size control which may occur in the small droplet regime, $R\ll\ell$. Suppose the reactions inside the droplet occur unhindered as in the dilute phase, meaning that, to leading order, $s_{\mathrm{l}}\simeq k_{\mathrm{s}\to\mathrm{ns}}$ (where we have taken $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\simeq 1$ and $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\simeq 0$). In that case, $R^{*}\sim(D_{\mathrm{g}}\Delta\varphi_{\mathrm{s}}/k_{\mathrm{s}\to\mathrm{ns}})^{1/2}\sim\ell\Delta\varphi_{\mathrm{s}}^{1/2}$, meaning that $R^{*}/\ell\ll 1$ and size control is bound to occur and thus the $R\gg\ell$ regime is inaccessible. To avoid this constraint, the reactions in the dense phase must occur with a reduced rate by excluding the enzymes or a fuel for reactions from the droplet. This can be achieved if, e.g., these catalyzers are bulky molecules and the droplets are very dense, or the sticky protein and the catalyzers have opposite hydrophilicities [15], or more generally if the sticky proteins inside the droplet apply a large entropic pressure against intruders [43]. We thus suppose that one of these mechanism causes the total reaction rate inside the droplet to be $s_{\mathrm{l}}<k_{\mathrm{s}\to\mathrm{ns}}\Delta\varphi_{\mathrm{s}}$, so $R$ may become comparable to or larger than $\ell$.
The second trouble may be the finite size of biological systems. For reasonable diffusivities $D_{\mathrm{g}}\sim 10^{-1}-10^{2}\mathrm{\mu m}^{2}\mathrm{sec}^{-1}$ and conversion rates $k_{i\to j}\sim 10^{-3}-10^{-1}\mathrm{sec}^{-1}$ [19], the distance between reactions may therefore vary between $\ell\sim 10^{-1}-10^{2}\mathrm{\mu m}$. The upper limit is larger than the typical eukaryotic cell size ($10\mathrm{\mu m}$), meaning that the accelerated ripening regime may not be attainable for slow reactions in biological systems. Therefore, we assume that the system size is sufficient to attain the $R\gtrsim\ell$ regime.
Once the regime $R>\ell$ is reached, we note that growth without ripening can be somewhat easily avoided. Since size control is overcome, $s_{\mathrm{l}}\sim D_{\mathrm{g}}\Delta\varphi_{\mathrm{s}}/\ell^{2}\sim(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\Delta\varphi_{\mathrm{s}}$, where $\Delta\varphi_{\mathrm{s}}\ll 1$. It is thus also possible to further require $s_{\mathrm{l}}\lesssim(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})\lambda/\ell$ for $\lambda/\ell\ll 1$. Under these conditions, accelerated ripening is achieved.
However, the third problem is how long does accelerated ripening continue? For our approach to hold in the absence of reactions (left panel of Fig. 1), we assumed small $\Delta\varphi_{\mathrm{s}}=\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$. To achieve meaningful acceleration, we need the deactivation reaction to be fast, $k_{\mathrm{s}\to\mathrm{ns}}\gg k_{\mathrm{ns}\to\mathrm{s}}$. This means, that in the active case, where $\Delta\varphi_{\mathrm{s}}=k_{\mathrm{ns}\to\mathrm{s}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$, supersaturation may not be large enough to produce an extended period of acceleration until ripening stops (let alone producing $\Delta\varphi_{\mathrm{s}}>0$ to allow ripening to occur in the first place). Moreover, the relevant supersaturation for the $R\gg\ell$ regime is in fact the even smaller $\Delta\varphi_{\mathrm{s}}^{\mathrm{eff}}$ of Eq. (28) (where $s_{\mathrm{l}}\sim k_{\mathrm{s}\to\mathrm{ns}}\Delta\varphi_{\mathrm{s}}$ with the above arguments). Therefore, not only must $s_{\mathrm{l}}$ be small, but $\phi_{\mathrm{s}\mathrm{g}}^{\circ}$ must be very small, too, so the solution remains effectively supersaturated.
In total, accelerated ripening can be realized in scenarios where there are almost no reactions inside the dense phase (so growth will not be hindered), the gas-phase reactions are not too slow (so they occur on lengthscales smaller than the system size), and the sticky species is almost insoluble (so the active solution would indeed be supersaturated).
Figure 4: Active acceleration of Ostwald ripening. The solid red curve is the theoretical prediction of the acceleration factor, $1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}$. The experimental data points are obtained from Ref. [15] in a case where the fuel concentration was kept constant up to an error. No fitting parameters were employed.
@@ IV.3 Realizability of accelerated ripening
An encouraging artificial chemical system that was shown to exhibit accelerated ripening according to this mechanism is reported in Ref. [15]. There, the sticky species is a strongly hydrophobic molecule which, upon exposure to water, breaks down to form a nonsticky (hydrophilic) molecule. Introducing a specialized fuel molecule allows conversion of the hydrophilic species through a sequence of reactions into the hydrophobic one. Since the sticky species creates a hydrophobic condensate impenetrable to water, its degradation does not occur at all inside droplet ($s_{\mathrm{l}}=0$). At the same time, since the degradation rate, $k_{\mathrm{s}\to\mathrm{ns}}$, is faster than the formation in the dilute phase, $k_{\mathrm{ns}\to\mathrm{s}}$, indeed an acceleration factor of up to a $100$-fold is reported [15].
@@
In Fig. 4 we plot our theoretical prediction along with the experimental results of Ref. [15], [^7] These three datapoints are obtained from experiments in which the fuel concentration was kept constant. Seven more datapoints are available in Ref. [15] from experiments where all of the fuel was introduced at the beginning, so during ripening only degradation of the hydrophobic species takes place. In that case, the accelerated ripening is enabled by accelerated shrinkage of small droplets. Our theory is quasistatic, and hence does not apply to this regime. with no fitting parameters. We estimate the reactions rates from measurements and numerical estimates available in the supplemental information of Ref. [15]. The acceleration factor is the ratio between slopes of the growth law of $R^{3}$ in the active and passive cases. Since the hydrophobic species does not exist without the fuel molecule, the growth rates for the passive cases are estimated via the Lifshitz-Slyozov law using measurements of surface tensions and solubilities. The agreement is not perfect, likely due to the approximations involved in estimating the passive growth rate, as well as $k_{\mathrm{ns}\to\mathrm{s}}$, and the protocol of periodic addition of the fuel that drives $k_{\mathrm{ns}\to\mathrm{s}}$. At the same time, the overall acceleration trend is reassuringly reproduced. Ref. [15] thus provides evidence that this singular means of accelerating ripening is indeed experimentally realizable.
V Discussion
@@
In this work, we theoretically studied the possibility of accelerating Ostwald ripening via chemical activity. We employed standard tools from linear-irreversible thermodynamics, Lishitz-Slyozov theory, and their extensions to include reactions to obtain explicit expressions for the diffusion-limited growth of droplets under ripening. We found that reactions in the sticky-particle-poor phase increase the flux of material out of small droplets and into large droplets. Indeed, if growth-hindering sticky-to-nonsticky reactions inside droplets can be avoided, we find acceleration of Ostwald ripening by an arbitrarily large factor, but with the same $R\sim t^{1/3}$ scaling imposed by particle conservation. Owing to the generality of our theory, we learn by elimination that this scenario in which reactions only occur in the dilute phase is indeed the only means to achieve accelerated Ostwald ripening. Our theoretical claims are qualitatively supported by recent experiments in an artificial chemical system [15]. Therefore, acceleration of ripening is another feat of active processes, joining the current ranks of size control [3, 4, 2, 5, 1], arrested ripening [6], Turing instabilities [7, 8], and bubbly phases [9].
The generality of our theory has been made possible through a series of simplifying assumption. Notably, we assumed that droplets are sparse, slowly growing, and that the source of activity does not modify surface thermodynamics. While these approximations are often reasonable, to actually achieve meaningful accelerated ripening is not trivial, requiring three components: (1) Reactions must be prevented inside the droplets. (2) The interconversion reactions must preferably form nonsticky species ($k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}>1$). (3) The overall concentration of material must be high enough that at chemical equilibrium in the dilute phase, the sticky species is supersaturated; otherwise droplets will not form. The experimental system of Ref. [15] follows all these assumptions to the letter, although the fact that, to our knowledge, it is the only instance where an acceleration by these means was obtained points to the challenge of realizing these conditions.
@@
Our theory has ignored droplet diffusion and coalescence — another means with which droplet growth can occur. Biomolecular condensates typically do not have a large surface tension, so growth by coalescence might in fact be the dominant effect driving the growth of natural biocondensates in cells [18]. To complicate matters further, growth by coalescense also exhibits $R\sim t^{1/3}$ scaling [25, 26], so distinguishing among growth by ripening from coalescence typically requires dynamic microscopy. Clearly, the interplay of reactions with droplet coalescence is a relevant future direction, as is the competition between the two growth mechanisms in active systems. We also ignored many other cellular complexities which interact with condensates [1]. In the context of the pyrenoid, these include membrane wetting mediated by molecular tethers [44] and the fact that two interacting components are required for droplet formation (Rubisco and EPYC1), rather than just a single “sticky” particle [23].
@@
On a theoretical level, our work is a manifestation of an overall theme concerning the ability to accelerate thermodynamic relaxation processes. One extensively studied example is the Mpemba effect, where an out-of-equilibrium system’s initial condition is chosen so it will undergo relaxation along a faster trajectory [45]. Other accelerated phenomena include shortcuts to adiabaticity [46, 47, 48], escape by active particles [49, 50], and search processes under stochastic resetting [51, 52, 53]. In this context, it is important to stress that not all types or degrees of acceleration are physically achievable. Activity typically breaks time-reversal symmetry, as in our case, but any accelerated process is still constrained by remaining unbroken conservation laws.
@@
On a practical level, to our knowledge, the acceleration method we report has only been realized in a chemical system [15]. Nevertheless, we posit that realizing accelerated Ostwald ripening in biocondensates is plausible. For example, advances in synthetic biology in principle allow the design of a phosphorylation-dependent switchable protein (like EPYC1) and a cognate kinase and phosphatase that are excluded from droplets [54], thus preventing conversion reactions in the dense phase. If rapid coarsening can be turned on or off at will, e.g., by optogenetics [55], the resulting size control of condensates would enable novel approaches to metabolic control [16, 17].
@@
To conclude, it is well established that living systems consume energy to carry out biological processes [56]. By driving molecular machinery “uphill” against chemical potential differences and thus breaking detailed balance, these active processes oppose relaxation to thermodynamic equilibrium. On the other hand, when relaxation to equilibrium is desired, it may be beneficial to accelerate the process. This fact is showcased by the ubiquity of enzymatic catalysis — a ubiquitous, non-energy-consuming means to accelerate biological reactions [57]. Accelerating processes such as phase separation as studied here but also membrane fission [58, 59] and fusion [60, 61, 62], DNA hybridization [63], and ligand-receptor binding [64] may be advantageous as well. It remains an exciting yet understudied question when and where cells use energy to accelerate thermodynamic relaxation.

### v2

@@ Abstract
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D_{\mathrm{eff}}}{R}\left[(\varphi_{\mathrm{eff}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})-\frac{\lambda_{\mathrm{s}\mathrm{g}}}{R}\right].$ (9)
Here, we identified the effective diffusion constant
$D_{\mathrm{eff}}=\left(1+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\right)D_{\mathrm{g}},$ (10)
which is enhanced by the dilute-phase reactions, and the effective supersaturation,
$\varphi_{\mathrm{eff}}^{\infty}=\frac{k_{\mathrm{ns}\to\mathrm{s}}\varphi^{\infty}-s_{\mathrm{l}}\ell_{\mathrm{l}}/\ell_{\mathrm{g}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}$ (11)
which is further reduced from the dilute-phase chemical-equilibrium value (Eq. (4)) due to the dense-phase reactions. Continuing from Eq. (9) along the lines of Lifshitz-Slyozov theory, we reproduce the same exact scaling but with an enhanced diffusivity, [^1] Note that due to the transient for $R\ll\ell$ and possibly altered nucleation rates [32, 1], the intercept $\langle R(0)\rangle^{3}_{\mathrm{a}}$ might differ from that of the passive case.
$\langle R(t)\rangle^{3}-\langle R(0)\rangle^{3}_{\mathrm{a}}=\frac{4}{9}D_{\mathrm{eff}}\lambda_{\mathrm{s}\mathrm{g}}t.$ (12)
This is our main result — activity enables Ostwald ripening to occur $(1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})$-fold faster compared to the passive case, Eq. (S100), if the introduced activity is such that dense-phase reactions are slow enough. Interestingly, the faster the dilute-phase switching from sticky to nonsticky compared to the reverse process, the faster is the ripening rate.
In addition to the restrictions on $s_{\mathrm{l}}$, there are two more limitations to this acceleration. First, the increase of $D_{\mathrm{eff}}$ occurs at the expense of having a lower sticky supersaturation, Eq. (4). Therefore, the saturation $\varphi^{\infty}$ must be sufficiently higher than $\phi_{\mathrm{s}\mathrm{g}}^{\circ}$ for the active solution to remain effectively supersaturated once the regime $R\gg\ell_{\mathrm{l}},\ell_{\mathrm{g}}$ is reached, i.e., $\varphi^{\infty}_{\mathrm{eff}}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}>0$. Second, biological systems are finite sized; e.g., a typical eukaryotic cell is $\sim 10\mathrm{\mu m}$. For reasonable protein diffusivities $D_{\mathrm{g}}\sim 10^{-1}-10^{2}\mathrm{\mu m}^{2}\mathrm{sec}^{-1}$ and conversion rates $k_{i\to j}\sim 10^{-3}-10^{-1}\mathrm{sec}^{-1}$ [19], the distance between reactions varies between $\ell_{\mathrm{l}},\ell_{\mathrm{g}}\sim 10^{-1}-10^{2}\mathrm{\mu m}$. Therefore, the system size must be sufficient to attain the $R\gg\ell_{\mathrm{l}},\ell_{\mathrm{g}}$ regime.
@@
It may not trivial to realize the above set of assumptions. Nonetheless, we argue for the feasibility of accelerating ripening by presenting recent experimental results as evidence for a workable protocol [15]. There, a strongly hydrophobic (sticky) molecule (so $\phi_{\mathrm{s}\mathrm{g}}^{\circ}\to 0$) converts to a hydrophilic (nonsticky) form upon exposure to water. Contrarily, a specialized fuel molecule converts the hydrophilic species through a sequence of reactions back into the hydrophobic from. Since the sticky species creates a hydrophobic condensate impenetrable to water, it is not lost by conversion at all inside droplets ($s_{\mathrm{l}}=0$). At the same time, in the dilute phase, the loss rate, $k_{\mathrm{s}\to\mathrm{ns}}$, is faster than the formation rate, $k_{\mathrm{ns}\to\mathrm{s}}$, so an acceleration factor of up to a $100$-fold is reported [15].
Figure 3: Active acceleration of Ostwald ripening. The solid red curve is the theoretical prediction of the acceleration factor — the ratio between the active ripening rate, Eq. (12), and the passive one, Eq. (S100), given by $D_{\mathrm{eff}}/D_{\mathrm{g}}=1+k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}}$. The experimental data points [15] correspond to a case where the fuel concentration was kept approximately constant. No fitting parameters were employed.
@@
In Fig. 3 we plot our theoretical prediction along with the experimental results [15], with no fitting parameters. [^2] These three datapoints are obtained from experiments in which the fuel concentration was kept constant. Seven more datapoints are available from experiments where all of the fuel was introduced at the beginning, [15] so during ripening only degradation of the hydrophobic species takes place. In that case, the accelerated ripening is enabled by accelerated shrinkage of small droplets. Our theory is quasistatic, and hence does not apply to this regime. The acceleration factor is the ratio between slopes of the growth law of $R^{3}$ in the active and passive cases. Since the hydrophobic species does not exist without the fuel molecule, the growth rates for the passive cases are estimated via the Lifshitz-Slyozov law using measurements of surface tensions and solubilities. The agreement is not perfect, likely due to the approximations involved in estimating the passive growth rate, as well as $k_{\mathrm{ns}\to\mathrm{s}}$, and the protocol of periodic addition of the fuel that drives $k_{\mathrm{ns}\to\mathrm{s}}$. At the same time, the overall acceleration trend is reassuringly reproduced. Thus, we are provided with evidence that this singular means of accelerating ripening is indeed experimentally realizable [15].
@@
To summarize, in this work, we theoretically studied the possibility of accelerating Ostwald ripening via chemical activity. We employed standard tools from reaction-diffusion and Lishitz-Slyozov theories to obtain explicit expressions for the diffusion-limited growth of active droplets under ripening. Indeed, if growth-hindering sticky-to-nonsticky reactions inside droplets can be avoided, we find acceleration of Ostwald ripening by an arbitrarily large factor, but with the same $R\sim t^{1/3}$ scaling imposed by particle conservation. Owing to the generality of our theory, we posit that this scenario in which reactions only occur in the dilute phase is indeed the only means to achieve accelerated Ostwald ripening. Our theoretical claims are supported by recent experiments in an organochemical solution [15].
@@
Our theory has ignored droplet diffusion and coalescence — another means with which droplet growth can occur. For instance, biomolecular condensates typically have a low surface tension, so growth by coalescence might in fact be the dominant effect driving their growth in cells [18]. To complicate matters further, growth by coalescense also exhibits $R\sim t^{1/3}$ scaling [33, 34], so distinguishing among growth by ripening from coalescence typically requires dynamic microscopy. Clearly, the interplay of reactions with droplet coalescence is a relevant future direction, as is the competition between the two growth mechanisms in active systems. We also ignored many other cellular complexities which interact with condensates [1]. In the context of the pyrenoid, these include membrane wetting mediated by molecular tethers [35] and the fact that two interacting components are required for droplet formation (Rubisco and EPYC1), rather than just a single “sticky” particle [23].
@@
On a theoretical level, our work is a manifestation of an overall theme in nonequilibrium statistical mechanics concerning the ability to accelerate thermodynamic relaxation processes. One extensively studied example is the Mpemba effect, where an out-of-equilibrium system’s initial condition is chosen so it will undergo relaxation along a faster trajectory [36]. Other accelerated phenomena include shortcuts to adiabaticity [37, 38, 39], escape by active particles [40, 41], and search processes under stochastic resetting [42, 43, 44]. In this context, it is important to stress that not all types or degrees of acceleration are physically achievable. Although activity typically breaks time-reversal symmetry (as in our case), any accelerated process is still constrained by remaining unbroken conservation laws.
@@
On a practical level, to our knowledge, the acceleration theory we report has only been achieved in a chemical system [15]. Nevertheless, we posit that realizing accelerated Ostwald ripening in biocondensates is plausible. For example, advances in synthetic biology allow the design of a phosphorylation-dependent switchable protein (like EPYC1) and a cognate kinase and phosphatase that are excluded from droplets [45], thus preventing conversion reactions in the dense phase. If rapid coarsening can be turned on or off at will, e.g., by optogenetics [46], the resulting size control of condensates would enable novel approaches to metabolic control [16, 17].
@@
To conclude, it is well established that living systems consume energy to carry out biological processes [47]. By driving molecular machinery “uphill” against chemical potential differences and thus breaking detailed balance, these active processes oppose relaxation to thermodynamic equilibrium. On the other hand, when relaxation to equilibrium is desired, it may be beneficial to accelerate the process. This fact is showcased by the ubiquity of enzymatic catalysis — a universal means to accelerate biological reactions [48]. Accelerating processes such as phase separation studied here but also membrane fission [49, 50] and fusion [51, 52, 53], DNA hybridization [54], and ligand-receptor binding [55] may be advantageous as well. It remains an exciting yet understudied question when and where cells harness energy to accelerate thermodynamic relaxation.